# In Defence of Post-hoc Explainability

### v1

@@ Abstract
The widespread adoption of machine learning in scientific research has created a fundamental tension between model opacity and scientific understanding. Whilst some advocate for intrinsically interpretable models, we introduce Computational Interpretabilism (CI) as a philosophical framework for post-hoc interpretability in scientific AI. Drawing parallels with human expertise, where post-hoc rationalisation coexists with reliable performance, CI establishes that scientific knowledge emerges through structured model interpretation when properly bounded by empirical validation. Through mediated understanding and bounded factivity, we demonstrate how post-hoc methods achieve epistemically justified insights without requiring complete mechanical transparency, resolving tensions between model complexity and scientific comprehension.
@@ 1 Introduction
The increasing adoption of machine learning (ML) in scientific research has created a fundamental tension between the opacity of complex ML models and the need for scientific understanding. Although ML models achieve unprecedented predictive performance across various scientific domains, from protein structure prediction to climate modelling, their complexity often renders them epistemically opaque, resistant to direct human understanding. This opacity presents a significant challenge to scientific practice, which traditionally relies on clear theoretical understanding and explanatory power.
The challenge has led to two dominant approaches in the ML community. One advocates for intrinsically interpretable models (Rudin, 2019), arguing that high-stakes (social) scientific applications demand transparent reasoning processes. The other develops post-hoc interpretability methods that attempt to explain already-trained complex models. This latter approach, while pragmatically valuable, has faced mounting scepticism: recent empirical studies have revealed limitations in current post-hoc methods (Hooker et al., 2019; Adebayo et al., 2022; Bilodeau et al., 2024), raising deeper philosophical questions about their epistemic justification.
In response to this tension, particularly regarding the epistemic status of post-hoc interpretability methods in scientific AI applications, we introduce Computational Interpretabilism (CI), a philosophical stance that acknowledges that while complete, factive explanations of complex AI systems might be currently unattainable or pragmatically limiting, we can achieve epistemically justified understanding within bounded approximations. This argument rests on two key principles: (1) scientific knowledge can emerge through structured interpretation of model behaviour, even without direct access to the model’s internal mechanisms; and (2) approximative explanations can provide epistemically justified scientific insights when we carefully define their limitations and verify their reliability through empirical testing.
@@ 2 On Interpretability
Human expertise has long been characterised by a distinctive gap between performance and "factive" explanation, particularly in domains where intuitive decision-making plays a central role (Dreyfus, 1972; Dreyfus and Dreyfus, 1986; Kahneman, 2003; Gobet and Chassy, 2008, 2009; Gobet, 2012). Experts across domains routinely make effective decisions while their subsequent explanations often involve post-hoc rationalisation rather than complete, factive accounts of their decision-making processes (Bilalić et al., 2008a, b).
This disjunction between performance and explanation is not a flaw but a fundamental feature of expertise, reflecting how complex knowledge is encoded and deployed in human cognition. Our acceptance of this epistemological limitation in human expertise, when contrasted with our scepticism toward similarly limited post-hoc explanations, raises fundamental questions about interpretability in scientific AI systems. Can explanation and performance be epistemically decoupled without compromising the justification of either? What can the limitations of human expertise teach us about epistemically justifying post-hoc interpretability methods? These questions challenge us to reconsider not just the technical aspects of post-hoc models, but the fundamental nature of explanation and understanding in scientific practice.
Before exploring these implications further, it is essential to establish key definitions and assumptions to frame our discussion. We begin by precisely defining what we mean by "interpretability".
@@
Interpretability in AI systems is fundamentally pluralistic (Zednik, 2021), encompassing multiple distinct concepts and serving diverse stakeholder needs. As Lipton (2016) and Beisbart and Räz (2022) emphasise, interpretability is not a monolithic concept but rather reflects different expectations, questions, and explanatory virtues depending on the stakeholder and context.
Different stakeholders approach interpretability with distinct questions and needs. Computer scientists typically focus on understanding how inputs are mechanistically processed to produce outputs – what we might call mechanistic interpretability (Nanda et al., 2023). In contrast, domain experts in scientific fields often seek to understand how model outputs inform or align with real-world phenomena – what we term phenomenological interpretability. While these perspectives are distinct, they are arguably reciprocal. As Ali et al. (2023) suggest, mechanistic interpretability often serves as a prerequisite for meaningful explanation, while the process of developing explanations can provide insights into the model’s functioning.
For scientific applications, which are the focus of this paper, interpretability takes on additional dimensions beyond just understanding model mechanics. It encompasses our epistemic capacity to understand and articulate how AI systems generate insights about natural phenomena in ways that advance scientific understanding. This broader conception is particularly important when considering post-hoc interpretability methods, whose criticisms often stem from evaluating them solely on their ability to faithfully explain a model’s learned function ($h^{*}(X)$) through some interpretable approximation ($p^{*}(X)$). However, in scientific contexts, the key relationship is not just between $h^{*}(X)$ and $p^{*}(X)$, but how both relate to the underlying natural phenomenon $f(X)$ being studied. (A detailed visualisation and explanation of these functions and their relationships is provided in Appendix A.) This three-way relationship suggests that post-hoc methods can serve a vital epistemological role: bridging between model computations and scientific understanding of natural phenomena, even when they may not perfectly capture all details of model behaviour. To further contextualise our discussion, we establish several key assumptions:
@@
3. Imperfect but Meaningful Approximations: We assume post-hoc methods provide approximations with imperfect but non-trivial fidelity to the original model. By this, we explicitly exclude methods that perform no better than random attribution (e.g., Hooker et al. (2019), Bilodeau et al. (2024)), focusing instead on approaches that, while not perfect, demonstrably capture meaningful patterns in model behaviour. While these approximations may not capture all nuances of model behaviour, they maintain sufficient accuracy to provide meaningful insights about both the model and the underlying phenomenon.
3 On Reliability and Justifiability
The parallel between human expertise and AI interpretability becomes particularly instructive when examining reliability and epistemic justifiability. Traditional epistemology distinguishes between two forms of justification (Pappas, 2005): internalist (requiring accessible reasons) and externalist (focusing on reliable processes). This framework illuminates how both human expertise and post-hoc AI interpretability can achieve epistemic justification despite limited explicit articulation.
Human experts, too, primarily provide externalist justification. Their expertise stems from extensive experience and training, which develops incomprehensible (yet sophisticated) cognitive processes connecting features of their current experience with vast stores of domain-specific knowledge (Dreyfus and Dreyfus, 1984, 1986; Kahneman and Klein, 2009; Gobet, 2012). Just as we trust our visual system without understanding its neural mechanisms, we accept expert judgment based on demonstrated reliability rather than complete explanation. Studies by Bilalić et al. (2008a, b) reveal that experts’ attention patterns often diverge from their reported reasoning, suggesting their explicit explanations are post-hoc reconstructions rather than precise accounts of their decision processes.
This recognition of post-hoc rationalisation in human expertise reshapes our perspective on AI interpretability: if we accept human expert judgments despite their post-hoc nature, we might similarly justify post-hoc AI interpretability methods when they demonstrate reliable knowledge generation. The key is not perfect mechanistic transparency but rather reliable processes for generating and validating insights.
Moreover, this parallel suggests that requiring complete internalist justification (full explicit explanation) from AI systems may be not just practically challenging but philosophically unnecessary (e.g., also see Sullivan (2022)). While intrinsic interpretability aligns with internalist approaches by emphasising direct access to reasoning processes, the success of human expertise suggests this may be unnecessarily restrictive. Just as human expertise explains implicit pattern recognition with explicit domain knowledge, AI systems might achieve scientific explanation through a combination of complex pattern recognition and post-hoc methods that reliably connect these patterns to domain knowledge.
This framework sets the stage for Computational Interpretabilism, which will show how post-hoc methods can achieve epistemic justification through bounded factivity and mediated understanding, even when their explanations are approximative. The key insight is that justification emerges not from complete mechanistic transparency but from reliable processes of knowledge generation and validation – a principle that applies equally to human expertise and AI systems.
4 Computational Interpretabilism
Computational Interpretabilism (CI) emerges as a philosophical framework that addresses one of the most pressing challenges in modern artificial intelligence: how to epistemically justify the use of post-hoc interpretability methods for scientific discovery. While machine learning models have demonstrated remarkable capabilities in analysing complex phenomena, their opacity raises fundamental questions about their role in generating scientific knowledge. CI provides a systematic approach to resolving this tension by establishing that post-hoc explanations can contribute legitimate scientific insights when they operate within specific epistemological boundaries (Beisbart and Räz, 2022). The framework synthesises multiple philosophical perspectives, including Sullivan’s link uncertainty, Andrews’ theory-laden understanding, and Freiesleben et al.’s holistic representationality, to demonstrate how approximative explanations can bridge the gap between model complexity and scientific understanding. A broader treatment of philosophical perspectives on AI interpretability, including discussions beyond these frameworks, is provided in Appendix C.
4.1 Philosophical Foundations
The epistemological relationship between machine learning models and scientific understanding emerges as a complex interplay of multiple philosophical challenges and frameworks, particularly in justifying post-hoc interpretability methods for scientific inquiry. Freiesleben et al.’s distinction between elementwise representationality (ER) and holistic representationality (HR) provides an essential foundation [^1] The key difference is that CI sees interpretability methods not just as tools for extraction but as active participants in knowledge creation. This leads to the concept of ”mediated understanding”. for understanding how modern ML models differ from traditional scientific models. This distinction gains deeper significance when viewed through multiple philosophical lenses: Andrews’ theory-ladenness, Beisbart and Räz’s factivity dilemma, Sullivan’s link uncertainty, and fundamental principles from philosophy of science such as epistemic accessibility (Longino, 1990; Kitcher, 2001) and intersubjective verifiability [^2] That scientific claims should be verifiable by multiple observers. (Kuhn, 1997; Popper, 2005). These perspectives together suggest that understanding emerges through mediated interpretation rather than direct access to model mechanics.
The shift from ER to HR that Freiesleben et al. propose can be understood as a practical response to the factivity dilemma – rather than attempting to maintain perfect fidelity at the component level (which would make models incomprehensible), HR acknowledges the necessity of holistic interpretation while providing rigorous methods for doing so through property descriptors. This approach aligns with Bilodeau et al.’s empirical findings that task-specific interpretability methods outperform general-purpose approaches, suggesting that effective scientific understanding requires targeted methods tailored to specific research contexts. However, as Andrews reminds us, these methods are inevitably theory-laden – the very choice of property descriptors and their implementation reflects our theoretical understanding and assumptions about both the phenomenon and the model. This theory-ladenness aligns with Douglas’s (2009) recognition that scientific inquiry is inherently value-laden, requiring careful consideration of how interpretability methods mediate between model behaviour and human scientific understanding.
Sullivan’s concept of link uncertainty provides a crucial bridge in understanding how post-hoc interpretability methods can be epistemically justified. These methods don’t eliminate link uncertainty, but rather help manage it by mediating between model behaviour and phenomenal understanding through rigorous, testable connections. This connects to Popper’s principle of falsifiability – interpretability methods must generate scrutinisable and potentially falsifiable claims about both model behaviour and phenomena. Beyond mere validation or falsification of existing knowledge, these methods can advance scientific understanding by revealing novel patterns and relationships that might not be apparent through traditional scientific approaches. The four-step framework proposed by Freiesleben et al. (formalisation, identification, estimation, and uncertainty quantification) can be seen as a systematic approach to reducing link uncertainty while acknowledging the theory-laden nature of scientific practice. Importantly, this framework demonstrates how post-hoc methods can be epistemically justified through their role in mediating between model behaviour, empirical validation, and scientific understanding.
The challenge of justifying post-hoc interpretability methods isn’t just a technical problem but a fundamental epistemological challenge that requires careful attention to how scientific understanding emerges through mediated interpretation. Drawing on pragmatist philosophy (Putnam, 1995), we must recognise that interpretability’s epistemic value lies in how it mediates between ML systems and human scientific understanding, facilitating oversight and integration with existing knowledge systems. Post-hoc interpretability methods emerge not just as technical tools for validation but as epistemological interfaces that actively participate in knowledge falsification and expansion, capable of generating new scientific insights while maintaining scientific rigour. They must make ML-generated knowledge accessible and comprehensible to the scientific community while enabling collaborative verification and critique, fundamentally shaping how we understand both ML systems and the phenomena they study.
Building on these philosophical foundations, CI establishes two key principles that together justify post-hoc interpretability methods in scientific ML. Mediated understanding reveals how scientific knowledge emerges through structured interactions between models, methods, and domain knowledge, while bounded factivity demonstrates why such mediated processes can be epistemically valid despite their approximative nature. Together, these principles provide a comprehensive philosophical framework that shows how post-hoc interpretability methods, when properly designed and validated, can contribute legitimately to scientific understanding.

### v2

@@ Abstract
This position paper defends post-hoc explainability methods as legitimate tools for scientific knowledge production in machine learning. Addressing criticism of these methods’ reliability and epistemic status, we develop a philosophical framework grounded in mediated understanding and bounded factivity. We argue that scientific insights can emerge through structured interpretation of model behaviour without requiring complete mechanistic transparency, provided explanations acknowledge their approximative nature and undergo rigorous empirical validation. Through analysis of recent biomedical ML applications, we demonstrate how post-hoc methods, when properly integrated into scientific practice, generate novel hypotheses and advance phenomenal understanding.
@@ 1 Introduction
Is explainable AI (XAI) in its peril (Weber et al., 2024)? Perhaps. However, this view has yet to gain mainstream acceptance. Driven by the increasing complexity and black-box nature of machine learning models, the field has witnessed an unprecedented surge in post-hoc explanation methods since 2017 (Saeed and Omlin, 2023; Nagahisarchoghaei et al., 2023; Retzlaff et al., 2024), with bibliometric analyses documenting exponential growth in XAI approaches (Sharma et al., 2024). This momentum manifests in substantial industry adoption exemplified by tools such as Captum (Kokhlikyan et al., 2020).
However, this popularity-driven surge cannot obscure the substantive criticisms that challenge the fundamental foundations of post-hoc explainability methods. Challenging both the theoretical premises and practical reliability of post-hoc explainability methods, these criticisms include:
  • Fidelity Guarantees: Critics such as Rudin (2019) argue that post-hoc explanation methods cannot provide genuine scientific understanding because they only approximate model behaviour rather than granting direct access to underlying mechanisms. This critique has gained formal computational support: recent complexity research demonstrates that generating explanations that are both interpretable and faithfully replicate model predictions is computationally intractable for most non-trivial models. Specifically, Bassan et al. (2024) and Adolfi et al. (2025) prove that local and global interpretability tasks, as well as discovering concise logical circuits for inner interpretability, are typically NP-hard or worse, precluding strong fidelity guarantees except in highly restricted cases.
  • Empirical Unreliability: A growing body of empirical investigations (Hooker et al., 2019; Laugel et al., 2019b, a; Slack et al., 2020; Adebayo et al., 2022; Bilodeau et al., 2024; Wei et al., 2024) has revealed pervasive limitations across the entire spectrum of post-hoc interpretability methods – spanning local and global approaches, model-agnostic and model-specific techniques alike. For instance, (Wei et al., 2024, pp. 7–10) addressed the issue of robustness in post-hoc methods, demonstrating that many popular methods fail to provide stable explanations. Most notably, Bilodeau et al. (2024) demonstrate that widely-adopted feature attribution methods (e.g., SHAP) can provably fail to outperform random guessing when inferring model behaviour in neural networks.
  • Scientific Methodology Concerns: Scientific domains remain particularly reluctant to accept post-hoc explanations as legitimate epistemic tools, with domain experts questioning their reliability for knowledge generation (Klein et al., 2024). Ghassemi et al. (2021) warn that post-hoc methods create ‘false hope’ in critical applications by appearing to provide understanding without reliable epistemic foundations. This critique fundamentally questions whether these methods can generate legitimate scientific knowledge rather than merely plausible-sounding narratives that satisfy human cognitive biases. (add Madsen et al. "empty argument")
  • Epistemological Imperative: Unlike other computational research domains that can rely on predictive accuracy or measurable outcomes for validation, interpretability researchers lack the empirical validation mechanisms that typically legitimate scientific claims – we rarely have the ground truth for benchmark evaluation in on the quality of generated explanations (Doshi-Velez and Kim, 2017; Yang et al., 2019), and the existence of a ground truth explanation is still highly debated (Bommer et al., 2024). This structural absence is compounded by the field’s methodological fragmentation, where no universal consensus exists on what methodology should be applied in evaluating XAI methods (Ehsan and Riedl, 2024), leading to repeated cycles where we take faithfulness for granted across different paradigms, later proving inadequate (Madsen et al., 2024).
Yet we argue that post-hoc interpretability methods can provide epistemically justified scientific understanding despite offering imperfect approximations of model behaviour. Rather than demanding complete, factive explanations – which may be unattainable for complex AI systems – we propose that scientific knowledge emerges through (1) structured interpretation that mediates between model behaviour and phenomenal understanding, and (2) approximative explanations that acknowledge their limitations whilst maintaining empirical accountability. This position recognises that scientific progress often proceeds through strategic simplifications, and post-hoc methods, when properly validated, can serve this same epistemic function.
The remaining paper is structured as follows: Section 2 examines the concept of interpretability in ML systems, contrasting intrinsic and post-hoc approaches. Through distinguishing between model understanding and phenomenon understanding, we establish the key theoretical relationships and assumptions that underpin our analysis. Section 3 introduces contemporary philosophical discussions relevant to ML interpretability, making these insights accessible to the broader ML community. We examine three interconnected themes: reliability and justifiability, scientific understanding and explanation, and scientific practice in the age of AI. Section 4 presents our integrated philosophical framework to defend post-hoc explainability. Through empirical case study, we also demonstrate how this framework translates theoretical principles into practical scientific discovery. Section 5 addresses and re-examines key criticisms of post-hoc methods. We offer constructive responses to concerns about approximation, fidelity, and confirmation bias. Section 6 concludes by outlining future research directions, identifying critical relationships in need of deeper theoretical examination and empirical validation. Throughout the paper, we minimise philosophical jargon where possible, with key concepts further explained in Appendix C for readers seeking additional clarification.
@@ 2 On Interpretability
Interpretability in ML systems is fundamentally pluralistic (Zednik, 2021), encompassing multiple distinct concepts and serving diverse stakeholder needs. As Lipton (2016) and Beisbart and Räz (2022) emphasise, the very notion of interpretability shifts and adapts based on stakeholder expectations, posed questions, and the explanatory virtues deemed most relevant within specific contexts.
This plurality in interpretation manifests most clearly in the distinct approaches taken by different stakeholders. Computer scientists typically gravitate towards what we might term model understanding – a focus on comprehending the properties of the model itself, including its input-output mappings, learned representations, and decision boundaries (e.g., seeking to understand model grokking through mechanistic interpretability (Nanda et al., 2023)). In contrast, domain experts in scientific fields often seek to understand how model outputs inform or align with real-world phenomena – what we might term phenomenon understanding. This distinction between model and phenomenon understanding is particularly salient when examining the scientific applications of ML interpretability in detail.
For scientific applications, which are the focus of this position, interpretability encompasses our epistemic capacity to understand and articulate how AI systems generate insights about natural phenomena in ways that advance scientific understanding. This relationship begins with a true function $f(X)$ representing the natural phenomenon of interest, where the model learns an approximation $h^{*}(X)$ within a hypothesis space $\mathcal{H}$ that represents the set of possible functions.
This broader conception is particularly important when considering post-hoc methods, whose criticisms often stem from evaluating them solely on their ability to faithfully explain a model’s learnt function ($h^{*}(X)$) through some interpretable approximation ($p^{*}(X)$). However, in scientific contexts, the key relationship is not just between $h^{*}(X)$ and $p^{*}(X)$, but how interpretation relates to the underlying natural phenomenon $f(X)$ being studied. This relationship manifests differently between intrinsically interpretable (1(a)) and post-hoc approaches (1(b)) (Appendix A. Whilst intrinsically interpretable models maintain direct human comprehensibility by constraining $h^{*}(X)$, post-hoc methods introduce an additional function $p^{*}(X)$ learnt from a space of possible explanations $\mathcal{P}$ to approximate $h^{*}(X)$’s behaviour in an interpretable way. These distinct approaches illustrate the fundamental tension between model capability and human comprehensibility in scientific ML applications, where post-hoc methods potentially offer the ability to capture more complex relationships between the model and the underlying phenomenon, albeit through an additional layer of interpretation.
These relationships encompass multiple interconnected aspects: the theoretical capacity of the hypothesis space to capture the true phenomenon ($f(X)$ and $\mathcal{H}$), the representational accuracy of training data ($f(X)$ and $X$), the model selection process ($\mathcal{H}$ and $h^{*}(X)$), the explanation generation mechanism ($\mathcal{P}$ and $p^{*}(X)$), the correspondence between model and explanation spaces ($\mathcal{H}$ and $\mathcal{P}$), and the properties of the post-hoc explanations themselves ($p^{*}(X)$).
However, this position paper will primarily address the relationship between phenomenon and explanation ($f(X)$ and $p^{*}(X)$). Whilst other relationships have been extensively studied within the ML community – such as $f(X)$ to $\mathcal{H}$ (e.g., Vapnik-Chervonenkis dimension suggesting that the capacity of the hypothesis space increases with both network width and depth (Bartlett and Maass, 2003)) and $\mathcal{H}$ to $h^{*}(X)$ (e.g., Structural Risk Minimisation demonstrating near-optimal learning guarantees (Stamouli et al., 2024)) – these discussions predominantly address engineering considerations rather than epistemological foundations.
This position aims to bring attention to the $f(X)$ to $p^{*}(X)$ relationship for three reasons. First, as AI models grow in complexity and opacity, $p^{*}(X)$ serves as our primary epistemic window into understanding complex phenomena $f(X)$ – a role that requires timely attention as model complexity grows. Second, this relationship bridges fundamental gaps in scientific AI, such as, between black-box models and human comprehension, between computational capabilities and scientific knowledge generation, and between theoretical frameworks and their practical applications in scientific discovery. And lastly, given the growing prevalence of post-hoc methods in studying complex natural phenomena, we require a philosophical framework to validate these methods as legitimate tools for scientific knowledge production.
To further contextualise our discussion, we establish several key assumptions, beyond the fundamental premise that $h^{*}(X)$ effectively approximates $f(X)$. These assumptions include:
@@
3. Imperfect but Justified Approximations: We assume that task-specific and properly validated post-hoc methods, when embedded within structured scientific practices, can provide empirically testable insights that advance scientific understanding. Whilst these methods may offer imperfect approximations of model behaviour, their epistemic value emerges not from perfect fidelity alone, but from their ability to generate falsifiable hypotheses, guide empirical investigations, and integrate with domain knowledge through systematic validation processes.
Having established these assumptions and motivations, we now turn to examine the philosophical foundations that underpin post-hoc explainability in scientific ML. This examination is particularly timely as the ML community increasingly grapples with questions that extend beyond purely technical considerations – questions that philosophy of science has wrestled with for centuries. By bridging these disciplines, we can leverage established philosophical frameworks to address emerging challenges in AI-driven scientific discovery, while simultaneously enriching both fields through their intersection. Moreover, as ML increasingly participate in scientific knowledge production, understanding their epistemological status becomes not merely an academic exercise, but a practical necessity for ensuring the validity and reliability of AI-generated scientific insights.
3 Philosophical Foundations
3.1 On Reliability and Justifiability
Traditional epistemology distinguishes between two forms of justification Pappas (2005): internalist approaches (which require that we can consciously access and articulate our reasons for believing something) and externalist approaches (which accept that reliable processes can justify beliefs even if we cannot fully explain how they work). Whilst intrinsic interpretability aligns with internalist approaches by emphasising direct access to reasoning processes, post-hoc methods align more naturally with externalist frameworks.
This distinction becomes particularly salient in the context of ML systems, where the inherent methodological opacity (inability to trace computational steps) and epistemic opacity (inability to understand what knowledge is encoded) of Deep Neural Networks (DNN) Humphreys (2009) challenges traditional notions of justification. This opacity manifests in two possible ways: through the algorithmic complexity that makes comprehensive understanding impossible, and through cognitive limitations that hinder human interpretation of the system’s behaviour Durán (2023).
Durán’s [2023] computational reliabilism (CR) addresses this challenge by proposing a framework that shifts focus from complete transparency to reliable belief-forming methods. CR delineates three categories of reliability indicators: (1) Technical Robustness of Algorithms, encompassing design, implementation, and maintenance factors; (2) Computer-based Scientific Practice, involving algorithmic implementation of scientific theories and expert assessment; and (3) Social Construction of Reliability, addressing the socially mediated processes of acceptance across scientific communities.
This framework has significant implications for post-hoc explainability methods in two ways. First, it suggests that scientific justification can emerge from the systematic validation of these methods through multiple reliability indicators, rather than requiring complete mechanistic understanding. Second, it provides a philosophical foundation for evaluating post-hoc methods based on their demonstrated reliability in connecting model behaviour to scientific phenomena, rather than their ability to provide complete transparency. This perspective reframes the epistemic status of post-hoc methods from approximate tools to potentially legitimate sources of scientific knowledge, provided they meet rigorous standards of reliability across technical, practical, and social dimensions.
3.2 On Scientific Understanding and Explanation
According to Sullivan (2022), the relationship between explanation and understanding in complex models hinges critically on the concept of “link uncertainty" – the gap between a model’s theoretical predictions and empirical reality. Sullivan argues that while models can be epistemically opaque (meaning their internal workings are not fully transparent), this opacity does not necessarily prevent them from providing genuine understanding, provided there is sufficient empirical evidence connecting the model to real-world phenomena. In other words, we do not necessarily need to fully understand how a model works internally; what matters more is understanding how the model connects to the real-world system it is studying.
Sullivan identifies three distinct types of explanatory questions we can ask about models: how the model itself works, how-possibly questions about potential mechanisms, and why/how-actually questions about real-world phenomena. The decisive factor in moving from how-possibly to how-actually explanations is reducing link uncertainty through scientific evidence. This evidence must connect the model’s theoretical insights to actual causal mechanisms in the target phenomenon. Importantly, Sullivan argues that understanding does not necessarily require complete knowledge of how a model works internally. Instead, what matters is the strength – whether it be the amount, kind or quality – of scientific and empirical evidence connecting the model’s predictions or insights to real-world phenomena.
This perspective gains additional significance through Beisbart and Räz’s [2022] analysis of the factivity dilemma – a fundamental tension between accuracy and comprehensibility in understanding DNNs. The principle of factivity (the requirement that explanations must be entirely true) demands that explanations be grounded in facts, yet modern DNNs have become so complex that we can only comprehend them through simplifications and idealisations. This creates an apparent paradox: explanations must either sacrifice accuracy for comprehensibility or maintain accuracy at the cost of being unusable.
This tension reflects a fundamental debate in philosophy of science about the relationship between explanation and understanding. Traditional accounts of scientific explanation, such as the Deductive-Nomological model, require factivity—that all elements of an explanation must be true. Yet scientific practice routinely employs strategic simplifications that sacrifice complete accuracy for tractable insights. This has led to two opposing views: factivists maintain that genuine understanding requires completely true explanations, treating simplifications as merely useful tools, whilst non-factivists argue that approximate models can provide legitimate understanding despite their imperfections. At the heart of this debate lies a crucial question: must understanding necessarily require truth, or can it emerge through useful approximations? When we refer to the “epistemic status” of post-hoc methods, we are asking precisely this – whether the knowledge claims they generate can be justified and relied upon, even when they offer only approximations rather than complete truth. Importantly, this suggests a distinction between interpretability, which aims for factual explanations of model behaviour, and scientific understanding of phenomena through post-hoc methods, which may not require the same level of factivity. This distinction offers a practical resolution: whilst complete explanations remain an important goal, we can achieve meaningful scientific understanding through carefully bounded approximations.
3.3 On Scientific Practice
The traditional view of ML as fundamentally distinct from conventional scientific methods – what scholars term the “distinctness claim" – has been challenged by Andrews’ [2023] analysis of theory-ladenness (how our theoretical assumptions inevitably shape what we observe and measure) in ML. The distinctness claim, suggesting that ML methods operate without prior theoretical assumptions, fundamentally misunderstands the nature of scientific data and practice. As Andrews demonstrates, even the most basic experimental designs reveal how data and scientific practice are inherently theory-laden, from the very act of investigation to the choice of measurement parameters.
Freiesleben et al. (2024) further advance this understanding by proposing “holistic representationality" (HR) – where we interpret the model’s overall behaviour – as an alternative to traditional “elementwise representationality" (ER) – where each model component directly represents something meaningful. Traditional scientific models followed what the authors call ER, where each model component – whether parameters, variables, or relationships – directly represented something meaningful about the phenomenon being studied. For instance, in a simple physics model, mass and velocity parameters directly correspond to physical properties. However, modern ML models, particularly DNNs, do not offer this kind of straightforward interpretation – their individual components (e.g., network weights) do not map clearly to real-world phenomena.
Rather than viewing this as a limitation, the authors proposed HR. Instead of trying to interpret individual components, they suggest analysing the model’s behaviour as a whole through what they call “property descriptors" (e.g., cPDP, cFI, SAGE, and PRIM for global property, while ICE, cSV, ICI and Counterfactuals for local property; see pp.21-25 for further details). This approach aligns with recent findings from Bilodeau et al. (2024), demonstrating that generic feature attribution methods can be unreliable for inferring model behaviour, but task-specific approaches can dramatically improve interpretability. While Freiesleben et al. provide a theoretical framework, Bilodeau et al. offer practical evidence of its importance, showing how domain-specific interpretability methods (e.g., perturbation) can be more reliable than general-purpose approaches like SHAP or Integrated Gradients.
This modern understanding reveals ML not as distant break from traditional scientific methods, but as a new set of tools requiring theoretical understanding and methodological rigour. The framework demonstrates how post-hoc methods, when properly constructed and validated, can serve as legitimate tools for scientific inquiry. This approach maintains scientific rigour whilst acknowledging the unique characteristics of modern ML systems, suggesting a path forward that neither overstates ML’s distinctness nor understates its methodological challenges.
4 Defending Post-hoc Explainability for Scientific ML
The philosophical foundations discussed above together suggest a more nuanced understanding of post-hoc interpretability methods in scientific ML. Building on these perspectives, we now develop a comprehensive framework that synthesises these insights and demonstrates their practical implications. We first present an integrated philosophical framework that establishes the epistemic validity (knowledge-generating legitimacy) of post-hoc methods in scientific inquiry, followed by a real-world case study that illustrates these principles in practice.
4.1 An Integrated Framework
The epistemological relationship between ML models and scientific understanding requires a synthesis of multiple philosophical frameworks to justify post-hoc interpretability methods in scientific inquiry. At its foundation lies Freiesleben et al.’s [2024] distinction between elementwise representationality (ER) and holistic representationality (HR). While traditional scientific models follow ER, where each model component directly represents a physical property, modern ML models necessitate HR, where understanding emerges from analysing system-level behaviours and patterns.
This shift from ER to HR gains deeper significance when considered alongside three complementary philosophical perspectives. Andrews’ [2023] theory-ladenness reveals how even our choice of interpretability methods reflects theoretical assumptions about both phenomena and models. Beisbart and Räz’s [2022] factivity dilemma highlights the inherent tension between accuracy and comprehensibility in complex models. Sullivan’s [2022] link uncertainty framework demonstrates how scientific understanding can emerge despite incomplete mechanistic knowledge.
The HR approach offers a practical resolution to these philosophical challenges. Rather than pursuing perfect component-level fidelity, it embraces holistic interpretation while maintaining scientific rigour. This aligns with empirical evidence that task-specific interpretability methods outperform general-purpose approaches Bilodeau et al. (2024), suggesting that scientific understanding requires carefully tailored methods for specific research contexts. Through this lens, post-hoc methods serve not to eliminate uncertainty, but to manage it through rigorous, testable connections between model behaviour and phenomenal understanding Popper (2005).
Drawing on pragmatist philosophy Putnam (1995), which values practical consequences over abstract truth, we propose that post-hoc methods’ epistemic value lies in their role as mediators between ML systems and human scientific understanding. This mediated nature of understanding, combined with the recognition that science often advances through strategic approximations, leads us to two key principles: mediated understanding and bounded factivity. Post-hoc interpretability methods thus serve as epistemological interfaces that actively participate in knowledge falsification and expansion, capable of generating new scientific insights when maintaining scientific rigour.